OpenStudy (anonymous):

Find the general form of the equation of the the circle. Center at the point (-4, -3); containing the point (-3, 3

OpenStudy (lalaly):

equation of circle \[(x-(-4))^2+(y-(-3))^2=r^2\]\[(x+4)^2+(y+3)^2=r^2\]

OpenStudy (lalaly):

tofind r substitute (-3,3) in x and y

OpenStudy (lalaly):

\[(-3+4)^2+(3+3)^2= 1+36=37\] so 37=r^2